DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, February 2, 2005 Padelford C-36 3:50 pm Degenerate second order operators and Sub-Riemannian geometry

Yakar Kannai (Weizmann Institute, Israel)

The geodesic flow on a Riemannian manifold plays an important role in the analysis of non-degenerate second order linear partial differential operators. For several degenerate operators Sub-Riemannian geodesics play an analogous role, but the structure is much more involved. Both real and complex approaches are considered and compared.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.